OpenStudy (megannicole51):

integrate ((y^6/3)+y^3-(y^3/3)-y^2))dy from 1 to 2 @agent0smith the answer is 87/14

OpenStudy (abb0t):

\[\int\limits \frac{y^6}{3}+y^3-\frac{y^3}{3}-y^2~dy\]

OpenStudy (anonymous):

youre soo pretty !!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (abb0t):

You think I don't know what i am?

OpenStudy (megannicole51):

thank you

OpenStudy (anonymous):

my pleasure!!!

OpenStudy (abb0t):

That's not the answer. I'm asking if that's the given integral.

OpenStudy (megannicole51):

yes it is

OpenStudy (megannicole51):

and yes it is

OpenStudy (abb0t):

\[\frac{1}{3}\int\limits y^6~dy + \int\limits y^3~dy - \frac{1}{3} \int\limits y^3~dy - \int\limits y^2~dy \] then, evaluate the integrals, as you should normally would. Then following the Fundamental Theorem of Calculus part II, where F(b)-F(a) Use that property to solve. Where \(\sf \color{red}{b = 2}\), \(\sf \color{limegreen}{a = 1}\)

OpenStudy (abb0t):

For instance, the first integral, you would get \(\sf \color{blue}{\frac{1}{21}y^7}\), now, you do \(\sf \color{orange}{\frac{1}{21}[(2)^7-(1)^7]=}\)....

OpenStudy (megannicole51):

okay i got it now thank you so much

OpenStudy (abb0t):

You can finish the rest now, right?

OpenStudy (megannicole51):

yup:)

OpenStudy (megannicole51):

actually hold on... so i got \[\frac{ y ^{7} }{ 21 }+\frac{ y ^{4} }{ 4 }-\frac{ y ^{4} }{ 12 }-\frac{ y ^{3} }{ 3 }\] is that right so far?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Yes.

OpenStudy (megannicole51):

jk i got it...it is right

OpenStudy (megannicole51):

then you just plug in:) thank you

OpenStudy (abb0t):

Now you need to apply fundamental theorem of calculus and solve.

OpenStudy (megannicole51):

yup got it:) thanks! sorry i doubted myself

OpenStudy (agent0smith):

Doubt yourself less, Meg :P

OpenStudy (megannicole51):

hehe yeah yeah!

OpenStudy (agent0smith):

You're very smart! :)

OpenStudy (abb0t):

@agent0smith she's like 12. Stop.

OpenStudy (agent0smith):

She's a university student.

OpenStudy (zarkon):

For all you know 'she' could be a 400lb university custodian. You can be what you want to be on the internet.

OpenStudy (abb0t):

Are you implying something there?

OpenStudy (zarkon):

I am an android. I don't imply anything.

OpenStudy (abb0t):

I'm on to you.

OpenStudy (anonymous):

what ??????

OpenStudy (zarkon):

don't tell Dr. Soong.

OpenStudy (anonymous):

umm

OpenStudy (megannicole51):

um excuse me but I'm 20 and that is my actual picture and i go to the university of colorado so no im not 12 or fat thank you very much

OpenStudy (zarkon):

Since no one lies on the internet I believe you ;)